Please show all the steps you've done to reach the final answer. I am trying to learn, so please show your work. Typing your answer is important.

1) A metal plating company is considering four different methods for recovering by-product heavy metals from a manufacturing site’s liquid waste. The investment costs and annual net incomes associated with each method have been estimated. All methods have an 8-year life; the MARR is 11% per year, and an AW-based ROR analysis is required.

(a) If the methods are independent (because they can be implemented at different plants), which ones are acceptable?

(b) If the methods are mutually exclusive, determine which one should be selected.

Method

First Cost, $

Salvage Value, $

Annual Net Income, $/Year

A

−30,000

+1000

+4000

B

−36,000

+2000

+5000

C

−41,000

+500

+8000

D

−53,000

−2000

+10,500

2) The estimated annual cash flows for a proposed municipal government project are costs of $710,000 per year, benefits of $890,000 per year, and disbenefits of $210,000 per year. Calculate the conventional B/C ratio at an interest rate of 5% per year, and determine if it is economically justified.
Answer #1
Year A B C D
0 -30000 -36000 -41000 -53000
1 4000 5000 8000 10500
2 4000 5000 8000 10500
3 4000 5000 8000 10500
4 4000 5000 8000 10500
5 4000 5000 8000 10500
6 4000 5000 8000 10500
7 4000 5000 8000 10500
8 5000 7000 8500 8500
IRR 2.1% 3.4% 11.3% 11.1%

Formula: IRR (values)

IRR for C and D > MARR , therefore they shall be chosen independently
Mutually exclusive, C shall be chosen with highest IRR

2) Conventional B/C ratio = (PV of benefits - PV of disbenefits) / PV of costs
let n be infinity, PV = Annual payment/ r
B/C = (890000-210000)/710000 = 0.96